# Large Language Models as Span Annotators

### v1

@@ G.3 D2T-Evalgsmarena domain
Llama 3.3 0.165 0.062 0.144 0.118 0.276 0.081 0.189 0.108 0.077 0.456
GPT-4o 0.127 0.068 0.169 0.078 0.195 0.072 0.181 0.109 0.062 0.453
Claude 3.7 0.214 0.146 0.290 0.121 0.240 0.133 0.263 0.130 0.140 0.656
DeepS. R1 0.223 0.183 0.373 0.059 0.121 0.090 0.183 0.093 0.160 0.686
o3-mini 0.172 0.201 0.429 0.078 0.167 0.113 0.240 0.127 0.192 0.689
Gem. 2-FT 0.234 0.179 0.352 0.098 0.193 0.127 0.249 0.122 0.161 0.675
@@ G.4 D2T-Evalopenweather domain
Llama 3.3 0.209 0.092 0.270 0.118 0.346 0.103 0.303 0.200 0.053 0.272
GPT-4o 0.404 0.164 0.298 0.158 0.288 0.161 0.293 0.132 0.122 0.377
Claude 3.7 0.550 0.246 0.370 0.291 0.437 0.266 0.401 0.135 0.189 0.429
DeepS. R1 0.454 0.215 0.391 0.129 0.234 0.162 0.293 0.131 0.124 0.533
o3-mini 0.552 0.317 0.437 0.259 0.357 0.285 0.393 0.108 0.259 0.481
Gem. 2-FT 0.458 0.197 0.366 0.241 0.447 0.217 0.403 0.186 0.155 0.368
@@
G.5 MT-Eval – Main Results
@@
Llama 3.3 0.182 0.048 0.085 0.228 0.376 0.071 0.124 0.053 0.015 0.392
GPT-4o 0.158 0.063 0.107 0.169 0.291 0.081 0.137 0.057 0.076 0.428
Claude 3.7 0.301 0.117 0.174 0.304 0.435 0.151 0.223 0.071 0.131 0.628
DeepS. R1 0.177 0.077 0.121 0.119 0.187 0.084 0.132 0.047 0.058 0.631
o3-mini 0.197 0.085 0.143 0.137 0.234 0.096 0.162 0.066 0.100 0.646
Gem. 2-FT 0.257 0.104 0.172 0.117 0.211 0.102 0.174 0.071 0.066 0.710
@@
G.6 MT-Eval – By Language
Language $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
@@ G.6 MT-Eval – By Language
G.7 Propaganda – Main Results
@@
Llama 3.3 0.336 0.069 0.253 0.063 0.231 0.066 0.242 0.176 0.092 0.343
GPT-4o 0.344 0.092 0.294 0.036 0.114 0.051 0.164 0.113 0.066 0.234
Claude 3.7 0.460 0.109 0.277 0.088 0.225 0.098 0.249 0.151 0.155 0.113
DeepS. R1 0.354 0.081 0.257 0.058 0.182 0.067 0.213 0.146 0.091 0.476
o3-mini 0.418 0.152 0.431 0.065 0.183 0.091 0.257 0.166 0.154 0.517
Gem. 2-FT 0.560 0.106 0.277 0.186 0.484 0.135 0.353 0.218 0.202 0.493
Table 20: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on Propaganda.
@@
Table 21: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on Propaganda. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
G.8 Confusion Matrices
@@ G.9 Manual Analysis
Appendix H Related Work (extended)
@@
As simple numeric scores can be subjective and unreliable, the field is shifting towards annotating erroneous spans. The MT community has adopted a detailed error taxonomy, known as MQM (Lommel et al., 2014; Mariana, 2014, Multidimensional Quality Metrics). In MQM, each erroneous span in the translation has an assigned category, such as accuracy/terminology and severity, such as minor/major/critical. The final translation score is defined as a weighted sum of the error spans based on the error categories and severities. However, this leads to many ties and is problematic when judging longer translations. In addition, this protocol requires a very expensive expert human labor. For these reasons, Kocmi et al. (2024b), based on prior work by Popovic (2020), introduce ESA (Error Span Annotation) for machine translation evaluation. This protocol drops the error categorization, which is primarily useful for diagnosis and not evaluation, simplifies severities, and has the annotators also assign the final score after they have done the span annotations. To make large-scale annotation more efficient, Zouhar et al. (2024) use a quality estimation system that pre-highlights potential error spans in the translation. The annotators then “post-edit” the error annotations and assign the final score.
@@
Relatively few works have addressed evaluation as error span annotation with LLMs. To obtain error spans for translation evaluation, GEMBA-MQM (Kocmi & Federmann, 2023) promps an LLM with annotation guidelines based on MQM evaluation protocol. InstructScore (Xu et al., 2023) and TigerScore (Jiang et al., 2024a) fine-tune smaller LLMs on synthetic data generated by GPT-4 to generate error spans. Although computationally more efficient than prompting larger LLMs, these two metrics either require reference texts, or are limited to predefined evaluation aspects. Kartáč et al. (2025) apply an ensemble of open-weight LLMs as evaluators to a variety of tasks and aspects, generating both numeric scores and error spans with detailed explanations. To address potential data contamination, Kasner & Dušek (2024) collect novel structured data and model outputs in five domains, and collect span annotations for semantic accuracy errors with both crowdworkers and GPT-4.

### v2

@@ E.3 D2T-Evalgsmarena domain
Llama 3.3 0.165 0.107 0.247 0.122 0.282 0.114 0.263 0.149 0.077 0.456
GPT-4o 0.127 0.107 0.269 0.082 0.205 0.093 0.233 0.140 0.062 0.453
Claude 3.7 0.214 0.176 0.348 0.152 0.300 0.163 0.322 0.159 0.140 0.656
DeepS. R1 0.223 0.199 0.405 0.089 0.180 0.123 0.249 0.126 0.160 0.686
o3-mini 0.172 0.225 0.473 0.115 0.241 0.152 0.320 0.168 0.192 0.689
Gem. 2-FT 0.234 0.218 0.426 0.128 0.249 0.161 0.314 0.153 0.161 0.675
@@ E.4 D2T-Evalopenweather domain
Llama 3.3 0.209 0.112 0.327 0.118 0.344 0.115 0.335 0.220 0.053 0.272
GPT-4o 0.404 0.196 0.354 0.160 0.289 0.176 0.318 0.142 0.122 0.377
Claude 3.7 0.550 0.274 0.411 0.292 0.437 0.283 0.424 0.141 0.189 0.429
DeepS. R1 0.454 0.241 0.436 0.143 0.258 0.180 0.324 0.144 0.124 0.533
o3-mini 0.552 0.362 0.496 0.267 0.366 0.307 0.422 0.115 0.259 0.481
Gem. 2-FT 0.458 0.226 0.415 0.249 0.458 0.237 0.435 0.198 0.155 0.368
@@
E.5 MT-Eval – Main Results
@@
Llama 3.3 0.182 0.121 0.200 0.229 0.378 0.155 0.257 0.102 0.014 0.392
GPT-4o 0.158 0.141 0.240 0.195 0.327 0.156 0.266 0.110 0.076 0.428
Claude 3.7 0.301 0.226 0.325 0.335 0.469 0.262 0.373 0.111 0.131 0.628
DeepS. R1 0.177 0.169 0.268 0.183 0.280 0.168 0.262 0.094 0.058 0.631
o3-mini 0.197 0.169 0.291 0.161 0.275 0.160 0.275 0.115 0.100 0.646
Gem. 2-FT 0.257 0.184 0.312 0.180 0.339 0.173 0.304 0.130 0.066 0.710
@@
E.6 MT-Eval – By Language
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
@@ E.6 MT-Eval – By Language
E.7 Propaganda – Main Results
@@
Llama 3.3 0.336 0.070 0.243 0.063 0.219 0.066 0.230 0.164 0.092 0.343
GPT-4o 0.344 0.095 0.293 0.038 0.115 0.054 0.166 0.112 0.066 0.234
Claude 3.7 0.460 0.110 0.274 0.096 0.239 0.103 0.255 0.152 0.155 0.113
DeepS. R1 0.354 0.083 0.246 0.062 0.182 0.071 0.209 0.138 0.091 0.476
o3-mini 0.418 0.152 0.411 0.066 0.179 0.092 0.249 0.157 0.154 0.517
Gem. 2-FT 0.560 0.106 0.268 0.190 0.477 0.136 0.343 0.207 0.202 0.493
Table 20: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
@@
Table 21: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the Propaganda. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.8 Confusion Matrices
@@ E.9 Manual Analysis
Appendix F Related Work (extended)
@@
As simple numeric scores can be subjective and unreliable, the field is shifting towards annotating erroneous spans. The MT community has adopted a detailed error taxonomy, known as MQM (Lommel et al., 2014; Mariana, 2014, Multidimensional Quality Metrics). In MQM, each erroneous span in the translation has an assigned category, such as accuracy/terminology and severity, such as minor/major/critical. The final translation score is defined as a weighted sum of the error spans based on the error categories and severities. However, this leads to many ties and is problematic when judging longer translations. In addition, this protocol requires a very expensive expert human labor. For these reasons, Kocmi et al. (2024c), based on prior work by Popovic (2020), introduce ESA (Error Span Annotation) for machine translation evaluation. This protocol drops the error categorization, which is primarily useful for diagnosis and not evaluation, simplifies severities, and has the annotators also assign the final score after they have done the span annotations. To make large-scale annotation more efficient, Zouhar et al. (2025) use a quality estimation system that pre-highlights potential error spans in the translation. The annotators then “post-edit” the error annotations and assign the final score.
@@
Relatively few works have addressed evaluation as error span annotation with LLMs. To obtain error spans for translation evaluation, GEMBA-MQM (Kocmi & Federmann, 2023) promps an LLM with annotation guidelines based on MQM evaluation protocol. InstructScore (Xu et al., 2023) and TigerScore (Jiang et al., 2024a) fine-tune smaller LLMs on synthetic data generated by GPT-4 to generate error spans. Although computationally more efficient than prompting larger LLMs, these two metrics either require reference texts, or are limited to predefined evaluation aspects. Kartác et al. (2025) apply an ensemble of open-weight LLMs as evaluators to a variety of tasks and aspects, generating both numeric scores and error spans with detailed explanations. To address potential data contamination, Kasner & Dušek (2024) collect novel structured data and model outputs in five domains, and collect span annotations for semantic accuracy errors with both crowdworkers and GPT-4.